OpenStudy (anonymous):

f g(x) = 3x + 5, what is g–1(x)?

OpenStudy (anonymous):

if not f*

OpenStudy (anonymous):

inverse of g?

OpenStudy (anonymous):

yeah g^-1

OpenStudy (anonymous):

you mean g^-1(x)

OpenStudy (anonymous):

sorry I typed this and then posted it without checking it.

OpenStudy (anonymous):

it is ok :)

OpenStudy (cwrw238):

g-1(x) is the inverse of g(x) g(x) = 3x + 5 find x in terms of g(x) 3x = g(x) - 5 can you now find x ? then replace g(x) by x and x by g(-1)x and thats your inverse

OpenStudy (anonymous):

it is 3x + 5=g(x)=y solve for x then switch variables (y and x)

OpenStudy (anonymous):

wait what? Im lost

OpenStudy (anonymous):

you take 3x+5=y and switch the variables? so its 3y+5=x?

OpenStudy (anonymous):

is it x+5/3?

OpenStudy (asnaseer):

you made a small eror in the sign of the result - please re-check your answer

OpenStudy (anonymous):

okay hold on

OpenStudy (asnaseer):

remember you started with: g(x) = y = 3x + 5 and then swapped x/y variables to get: x = 3y + 5

OpenStudy (anonymous):

ok...

OpenStudy (asnaseer):

so now you need to re-arrange the last expression so that you end up with "y = ..." do you know how to do that?

OpenStudy (anonymous):

subtract 3y from both sides

OpenStudy (anonymous):

x-3y=5 then you subtract x from both sides. so -3y=-x-5

OpenStudy (asnaseer):

well first we can just swap the left and right hand sides to get: 3y + 5 = x